OpenStudy (anonymous):

Let f(0) not equal to 0 and for x and y belong to R, 2f(x)f(y)= f(x+y)+f(x-y) is always true, then, express f(2x) in terms of f(x).

OpenStudy (anonymous):

What did you try?

OpenStudy (anonymous):

confused.

OpenStudy (anonymous):

How about this, could you express f(2t) in terms of f(t)?

OpenStudy (anonymous):

If you could find the f(2t) in the equation, then you probably can do this problem.

hartnn (hartnn):

put x=y

OpenStudy (anonymous):

yeap. I got it. let x=y so 2*f^2(x)=f(2x)+k because f(0)=k so f(2x)=2f^2(x)-k

hartnn (hartnn):

seems correct.